Can not figure out the answers to combine like terms.
1.) -3+13w -7w-5
2.) 5x^4+3x^2- 6x^4+x
3.) 2- 5+3y+5y^2+4y
4.) 2d^3-4d+6
Please help! :) Thanks


Answer by stanbon(75887)

combine like terms.
1.) -3+13w -7w-5 = 6w-8
--------------------------------
2.) 5x^4+3x^2- 6x^4+x = -x^4+3x^2+x
----------------------------------------
3.) 2- 5+3y+5y^2+4y = 5y^2+7y-3
----------------------------------------
4.) 2d^3-4d+6 = 2d^3 - 4d + 6
-----------------------------------
Cheers,
Stan H.